# Magnetic signals from oceanic tides: new satellite observations and applications

### v1

@@ 4.2 Conductivity of the oceanic mantle
We recall that numerical simulations used for the comparisons with observed ocean tidal magnetic signals were calculated using a global radial conductivity obtained by inversion of the new CI9 $M_{2}$ tidal signal (see Section 2.3). Since all constituents sense the same subsurface conductivity, they should tend towards these numerical simulations. To obtain further insights into how satellite-detected ocean tidal magnetic signals for each tidal constituent can sense the lithosphere-asthenosphere system individually, we designed the following parameter study: A 3-D conductivity model of the world ocean and marine sediments as described in Section 2.3 was used. Below these layers, a model with a 80-km thick layer followed by a half-space was adopted. In reality, the depth to the lithosphere-asthenosphere boundary (LAB) varies with location, but to first order it is controlled by the age of the oceanic crust [54] with an average depth of $\approx 70-80$ km [25].
@@
For every combination of the lithosphere and asthenosphere electrical conductivity, the normalized RMS error between observed and simulated fields at satellite altitude (430 km) was calculated. Standard deviation values for each constituent model from Figure 4) were used for normalization. The distribution of RMS error values for all tested combinations of the lithosphere-asthenopshere conductivity contrast are presented in Figure 7. Remarkably, all tidal constituents are sensitive to the conductivity contrast at the LAB. The steepest slopes around the minimum are seen for the strongest $M_{2}$ constituent, which has the best SN ratio, whereas the opposite is true for the smallest $Q_{1}$ constituent, although the global minimum is nevertheless clearly visible. Another noteworthy observation is that diurnal tides favour a more electrically conductive asthenosphere, which can be explained since electromagnetic fields at longer periods penetrate deeper and sense the generally more conductive deeper upper mantle [54]. The differences in the retrieved minimum can be explained by different sensitivity patterns, especially between semi-diurnal and diurnal components, as well as by the presence of noise. Additionally, these plots clearly reveal the non-linear nature of the inverse problem and the presence of several local valleys and ridges in the misfit function landscape, motivating the use of advanced non-linear optimization techniques for electromagnetic induction inverse problems.
@@ 5 Discussion and Conclusions
The detection of the tiny signals from the $Q_{1}$ tidal harmonic and its sensitivity to a plausible range of electrical conductivities in the upper mantle demonstrates the impressive accuracy of the satellite magnetometers and confirms low calibration errors that can be averaged out very well, at least at the considered frequencies. In this study we were however not able to reliably characterise other tidal constituents, such as $K_{1}$ or $S_{2}$. Despite their overall amplitudes being higher than that of $Q_{1}$ or even $O_{1}$ (Table 1), their periods (nearly) coincide with the solar daily harmonics. The magnetic field at these periods is dominated by ionospheric fields (see also the discussion in the introduction) even during the quiet conditions considered here. Therefore, further improvements in mitigating the effect of the ionosphere and high latitude currents are needed for progress on these constituents.
The model parameterisation given by eq. (3) targets the stationary harmonic magnetic signal due to a specific tidal constituent of interest. In other word, retrieved signals are independent of variability caused by factors such as seasonal [55] or long-term changes in the ocean climatology [56]. This approach is preferred for mantle induction applications because the electrical conductivity of seabed and mantle remain essentially constant over the two decades considered here. Since the dependency of tidal magnetic signals on the conductivity of the solid Earth is larger than the component caused by long-term ocean climatology changes, it is important to first characterise the solid Earth part of the signals before trying to retrieve and model putative climatology-driven effects. This appears to be the most consistent way of addressing this non-unique problem. The experience gained with extraction of stationary harmonic tidal magnetic signals will certainly be vital when attacking the problem of detecting oceanic magnetic signals due to other forcing mechanisms and detecting climatology-related changes.
@@ Acknowledgement
AG is grateful to Gary Egbert for insightful discussions on oceanic tides. We thank Dr. J. Baerenzung for providing us with the updated KALMAG models. We used the ChaosMagPy package [57] for some post-processing and plotting.

### v4

@@ 4.2 Conductivity of the oceanic mantle
We recall that numerical simulations used for the comparisons with observed ocean tidal magnetic signals were calculated using a global radial conductivity obtained by inversion of the new CI9 $M_{2}$ tidal signal (see Section 2.2). Since all constituents sense the same subsurface conductivity, they should tend towards these numerical simulations. To obtain further insights into how satellite-detected ocean tidal magnetic signals for each tidal constituent can sense the lithosphere-asthenosphere system individually, we designed the following parameter study: A 3-D conductivity model of the world ocean and marine sediments as described in Section 2.2 was used. Below these layers, a model with a 80-km thick layer of conductivity $\sigma_{\mathrm{L}}$ followed by a half-space of conductivity $\sigma_{\mathrm{A}}$ was adopted. In reality, the depth to the lithosphere-asthenosphere boundary (LAB) varies with location, but to first order it is controlled by the age of the oceanic crust [63] with an average depth of $\approx 70-80$ km [29].
@@
The lithosphere and asthenosphere electrical conductivities $\sigma_{\mathrm{L}}$ and $\sigma_{\mathrm{A}}$ were varied in the range $10^{-5}$ to $10^{1}$ S/m. For every combination of these values, the normalized RMS error between observed and simulated fields at satellite altitude (430 km) was calculated. Standard deviation values for each constituent model from Figure 4) were used for normalization. The distribution of RMS error values for all tested combinations of the lithosphere-asthenosphere conductivity contrast are presented in Figure 7. Remarkably, all tidal constituents are sensitive to the conductivity gradient across the LAB as is evident from the changes in the normalized RMSE upon variations in the conductivity of the lithosphere against that of the asthenosphere. The steepest slopes in the normalized RMSE around the minimum are seen for the strongest $M_{2}$ constituent, which has the best signal-to-noise ratio, whereas the opposite is true for the smallest $Q_{1}$ constituent, although the global minimum is nevertheless clearly visible. Another noteworthy observation is that diurnal tides favour a more electrically conductive asthenosphere, which can be explained since electromagnetic fields at longer periods penetrate deeper and sense the generally more conductive deeper upper mantle [63]. The differences in the retrieved minimum can be explained by different sensitivity patterns, especially between semi-diurnal and diurnal components, as well as by the presence of noise. Additionally, these plots clearly reveal the non-linear nature of the inverse problem and the presence of several local valleys and ridges in the misfit function landscape, motivating the use of advanced non-linear optimization techniques for electromagnetic induction inverse problems.
@@ 5 Discussion and Conclusions
The detection of the tiny signals from the $Q_{1}$ tidal harmonic and its sensitivity to a plausible range of electrical conductivities in the upper mantle demonstrates the impressive accuracy of the satellite magnetometers and confirms low calibration errors that can be averaged out very well, at least at the considered frequencies. In this study we were however not able to reliably characterise other tidal constituents, such as $K_{1}$ or $S_{2}$. Despite their overall amplitudes being higher than that of $Q_{1}$ or even $O_{1}$ (Table 1), their periods (nearly) coincide with the solar daily harmonics. The magnetic field at these periods is dominated by ionospheric fields (see also the discussion in the introduction) even during the quiet conditions considered here [11]. It seems justified to extrapolate this argument to ocean tides of lunar and solar origins in general. Specifically, the current trend of improving retrievals of the ocean tidal magnetic signals due to lunar constituents is likely to continue with ongoing high-quality space observations. In contrast, reliable separation of tidal ocean magnetic signals due to solar constituents (or constituents with periods that are very close to the daily period and sub-harmonics) depends largely on how well we can characterize and mitigate the effect of the ionosphere and high latitude currents. The latter challenge is beyond the realm of this study, and will require dedicated efforts in order for progress to be made [64].
The model parameterisation given by eq. (3) targets the stationary harmonic magnetic signal due to a specific tidal constituent of interest. In other word, retrieved signals are independent of variability caused by factors such as seasonal [46] or long-term changes in the ocean climatology [65]. This approach is preferred for mantle induction applications because the electrical conductivity of seabed and mantle remain essentially constant over the two decades considered here. Since the dependency of tidal magnetic signals on the conductivity of the solid Earth is larger than the component caused by long-term ocean climatology changes, it is important to first characterise the solid Earth part of the signals before trying to retrieve and model putative climatology-driven effects. This appears to be the most consistent way of addressing this non-unique problem. The experience gained with extraction of purely harmonic tidal magnetic signals will certainly be vital when attacking the problem of detecting oceanic magnetic signals due to other forcing mechanisms and detecting climatology-related changes.
@@
Data accessibility
All final models for the four considered tidal constituents, numerical simulation results and scripts to reproduce the figures are archived in a public repository https://github.com/agrayver/magtide. Swarm data is available at http://swarm-diss.eo.esa.int. CHAMP mission data is archived at dataservices.gfz-potsdam.de. Models such as the CHAOS, IGRF, ocean and sediments conductivity are available via cited references. The TPXO-9 model was accessed through the www.tpxo.net/global/tpxo9-atlas portal.
@@ Acknowledgement
AG is grateful to Gary Egbert for insightful discussions on oceanic tides. We thank Dr. J. Baerenzung for providing us with the updated KALMAG models. We used the ChaosMagPy package [66] for some post-processing and plotting. This work was supported by the Heisenberg Grant from the German Research Foundation, Deutsche Forschungsgemeinschaft (Project No. 465486300) and ESA Swarm DISC project No. 4000109587.